2. The resistance of a certain wire is 10 Ω. a) What would the resistance of the same wire be if it were twice as long? b) If it had twice the diameter?
1
Expert's answer
2020-04-09T09:25:42-0400

a)


"R'=\rho\frac{2l}{0.25\pi d^2}=2R=20\Omega"

b)


"R''=\rho\frac{l}{0.25\pi (2d)^2}=0.25R=2.5\Omega"